Amy bought some strings of beads. The number of strings was 16 more than the number of beads on each string. There were 57 beads in all. How many beads were on each string?
1
Expert's answer
2020-02-02T16:31:40-0500

Let x be the number of strings bought.

Let y be the number of beads in each string

Then,

x=16+y……….. (i)

xy=57……….(ii)

from equation (ii),x=57/y. When we substitute this value into equation (i), it becomes,

57/y=16+y

Multiply both sides by y,

57=16y+y2.

Rearrange the equation in the quadratic form

y2+16y-57=0

y=-19 or y=3

Pick the positive number y=3.

Hence there were 3 beads on each string.